A wallet contains a Php 20.00 bill, a Php 50.00 bill, and a Php 100.00 bill. A bill is taken out from the wallet and then replaced with a bill of the same amount that was taken away. If a bill is taken out from the wallet, what is the probability that the sum of the bills is

A. Php 70.00

B. Not Php 70.00

C. More than Php 40.00


1
Expert's answer
2021-03-18T03:58:35-0400

According to the question:

The wallet contains 1 20 Php bill

1 50 Php bill

1 100 Php bill


(A). The amount is taken with replacement

So, the probability of getting 70 Php bill = "2\times {1 \over 3 }\times {1 \over 3}={2\over 9}"

(B). P(Not 70)= "1 - \dfrac{2}{9}= \dfrac{7}{9}"

(C). P(> 40)= P(70) + P(120) + P(150)= "2\times ({1 \over 3 }\times {1 \over 3})+2\times({1 \over 3 }\times {1 \over 3})+2\times({1 \over 3 }\times {1 \over 3})= \dfrac{2}{3}"